OpenStudy (anonymous):

The graph of the solution to the inequality 1/3x-<5 has a closed circle at 2 and is shaded to the right @ranga can you help me? true or flase

OpenStudy (anonymous):

Is it false, but the circle would be closed?

OpenStudy (anonymous):

*because

OpenStudy (ranga):

Is the inequality:\[\Large \frac{ 1 }{ 3 }x \le 5\]?

OpenStudy (ranga):

It is not clear what the inequality is in the problem. Does it look like my previous reply?

OpenStudy (anonymous):

yes, but the < is not underlined

OpenStudy (ranga):

1/3x-<5 What is that between x and <

OpenStudy (anonymous):

1/3x-1<5

OpenStudy (anonymous):

sorry about that

OpenStudy (ranga):

1/3x - 1 < 5 add 1 to both sides 1/3x < 6 multiply both sides by 3 x < 18 which has nothing to do with closed circle at 2. So it is false.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

can you check a problem for me?

OpenStudy (ranga):

go ahead.

OpenStudy (anonymous):

-2x+3y=7 +2x +2x 3y/3= 7+2x y=(7+2x)/3 ??

OpenStudy (ranga):

y = 2/3x + 7/3

OpenStudy (ranga):

Were you asked to put the equation in slope intercept form?

OpenStudy (anonymous):

Solve the equation for the letter y

OpenStudy (ranga):

okay, if all it says is solve for y, you can stop at what you had at y=(7+2x)/3. Either one is fine.

OpenStudy (ranga):

I just split (7 + 2x) / 3 as 7/3 + 2x/3. But you can leave it either way and both are fine.

OpenStudy (anonymous):

okay, thank you so much!

OpenStudy (ranga):

yw.